the sum of the ages of carlos and rosa in 14 years will be equal to 2 times their present age. if at present age carlos is 8 years elder to rosa then what is Carlo's present age?

Answer by ikleyn(52787)

.
the sum of the ages of Carlos and Rosa in 14 years will be equal to 2 times their present age.
if at present age Carlos is 8 years elder to Rosa then what is Carlo's present age?
~~~~~~~~~~~~~~~

One equation is

    (C+14) + (R+14) = 2*(C+R),


or, equivalently,

    C + R = 28.    (1)


Another equation is 

    C - R =  8     (2)


Add equations (1) and (2) and get

    2C = 28 + 8 = 36;  C = 36/2 = 18.


So, Carl is 18 years old.    ANSWER

Solved.